OpenStudy (anonymous):

Please Help! Will Medal! Samuel took out a personal loan of 13,000 at 7% interest, how much interest will he pay after 4 years. Suppose he pays off the loan in 2 years instead of 4 how much money will he save in interest

OpenStudy (kittiwitti1):

I'm assuming this is an appropriate/sufficient answer: http://openstudy.com/study#/updates/54c93106e4b0331aa6852d42

OpenStudy (kittiwitti1):

I'll calculate to make sure, hold on

OpenStudy (anonymous):

The second part is the part I am struggling with, thank you though.

OpenStudy (kittiwitti1):

Well you just halve the years, no?

OpenStudy (kittiwitti1):

You can always check with a interest calculator - http://www.calculatorsoup.com/calculators/financial/simple-interest-plus-principal-calculator.php

OpenStudy (anonymous):

This is my equation I have gotten so far... 13,000 x 0.07 x4 = 3,640

OpenStudy (kittiwitti1):

Here: http://prntscr.com/6vlarx

OpenStudy (kittiwitti1):

So then: "First, converting R percent to r a decimal r = R/100 = 7%/100 = 0.07 per year, then, solving our equation" A = 13000(1 + (0.07 × 2)) =

OpenStudy (kittiwitti1):

Is that understandable? :)

OpenStudy (anonymous):

Yes, Thank you very much!

OpenStudy (kittiwitti1):

You're welcome ^^, please don't forget to close the question! Thanks for medal :)

OpenStudy (anonymous):

Ok, No problem :)